# Elastocapillarity-driven surface growth in tumour spheroids

### v1

@@
Elastocapillarity-driven surface growth
@@
D. Riccobelli email: davide.riccobelli@polimi.it
(MOX – Dipartimento di Matematica, Politecnico di Milano, Piazza Leonardo da Vinci 32, 20133, Milano, Italy)
@@ 1 Introduction
Cancer research frequently relies on surrogates to investigate the dynamics of tumour growth, ranging from experiments performed in vivo to in vitro. In the first type of experiments, tumours are implanted and grown in animals, such as mice [Stylianopoulos_2012]. Conversely, in vitro experiments involve the cultivation of tumour cells outside a living organism in the laboratory [Zanoni_2020]. A notable example is multicellular tumour spheroids, which are three-dimensional structures that self-assemble, forming spherical aggregates.
@@
In order to understand how boundary loads drive the inner mechanical state, a rheological description of spheroids is essential. Increasing evidence suggests that spheroids exhibit viscoelastic behaviour [Ambrosi_2017]. For example, similarly to fully developed tumours, spheroids tend to open when subjected to a radial cut [Colin_2018, Guillaume_2019, Kosheleva_2023]. If the spheroid obeyed a fluid-like behaviour governed by surface tension, it would reorganize to minimize surface area, returning to a spherical shape. Conversely, radially cut tumour spheroids display a heart-like shape, see Fig. 1. Moreover, even intact tumour spheroids can exhibit both spherical and cylindrical shapes depending on the initial number of cells [Guillaume_2019], suggesting that while surface tension plays a role in spheroid physics, other forces also influence its shape.
Figure 1: (Left) Sections of tumour spheroids with varying radii. Recently dividing cells are coloured green, while dying cells are indicated in blue (adapted from [Browning_2021]). (Right) An incised tumour spheroid two days post-seeding, with an initial cell count of 5000 (adapted from [Guillaume_2019]).
@@
During the initial days, only a thin outer layer of cells experiences circumferential tension, while the core remains in an isotropic stress state [Lee_2019]. This tensional “skin” at the spheroid’s boundary can be understood as an effect of tissue surface tension on the cell aggregate generated by intercellular cohesion forces and active contractility, as discussed in [Riccobelli_2020]. This surface phenomenon primarily arises from intercellular cohesion forces and has been shown to significantly influence the mechanical behaviour of cell aggregates [manning2010coaction, Riccobelli_2020], particularly at small scales such as those of a spheroid [Kosheleva_2023, Ang_2024].
The aim of this work is to construct a mathematical model of tumour spheroid growth able to explain the tumour opening following a radial cut. In Section 2, we propose the mathematical model of the spheroid, where tumour proliferation is described as a surface accretion mediated by surface tension. The study of spheroid incision is performed in Section 3, where we propose both a simplified analytical prediction of tumour opening and a quantitative analysis through numerical simulations and comparison with experimental results. Finally, the main results are summarized in Section 4, together with some concluding remarks.
2 Modelling of the elastocapillarity of tumour spheroids
@@
2.1 A continuum model of tumour spheroid with tissue surface tension
@@
In order to take into account the large elastic deformations of the tumour, we assume that the cellular aggregate can be described as a hyperelastic material. Let $\psi_{r}=\psi_{r}(\mathbf{F})$ be its strain energy density. The bulk elastic energy of the body is thus given by
@@
Since cellular aggregates are mainly composed of water, we can approximate their behaviour as incompressible, namely the deformation must obey the constraint
@@
where $\mu$ is the shear modulus, while :::: denotes the Frobenius scalar product of tensors, i.e. $\mathbf{A}:\mathbf{B}=\operatorname{tr}\mathbf{A}^{T}\mathbf{B}$.
2.2 Surface accretion and surface tension
While the model proposed in the previous section can describe the pure elastic behaviour of a spheroid, it is insufficient to capture what occurs as the spheroid grows. Indeed, spheroids left to grow for several days in an environment with large availability of nutrients can generate mechanical stress as a consequence of their growth. The increase in the opening length of cut spheroids grown for a few days [Guillaume_2019] suggests that the growing process enhances the natural opening induced by surface tension by generating a tensile hoop stress in the outer part of the spheroid.
@@ 2.2 Surface accretion and surface tension
Differently from previous models [Mascheroni_2016, Ambrosi_2017, Dolega_2021, Erlich_2023, Walker_2023], here we describe spheroid’s growth as a surface accretion, where new mass is added at the boundary rather than in the bulk. We start by considering a time-dependant reference configuration, i.e. $\Omega_{0}(t)$. Let $R_{t}$ be the radius of the (intact) tumour spheroid at time $t$, so that
@@
Such an operation is called material transplant. Generally speaking, the reference configuration may not be stress-free if the transplant map $\mathbf{P}(\boldsymbol{X})$ is not a gradient of a deformation function $\boldsymbol{\varphi}_{\mathbf{P}}(\boldsymbol{X})$, for a detailed discussion see [Epstein_2012, Goriely_2017]. In the following, the explicit dependence of $\mathbf{P}$ on the point is omitted unless necessary.
@@
$\begin{gathered}\Sigma_{RR}=\left\{\begin{aligned} &4\gamma\left(\frac{1}{R_{T}}-\frac{1}{R}\right),&&\text{if }R\geq R_{0},\\ &4\gamma\left(\frac{1}{R_{T}}-\frac{1}{R_{0}}\right),&&\text{if }R<R_{0},\end{aligned}\right.\\ \Sigma_{\Theta\Theta}=\left\{\begin{aligned} &2\gamma\left(\frac{2}{R_{T}}-\frac{1}{R}\right),&&\text{if }R\geq R_{0},\\ &4\gamma\left(\frac{1}{R_{T}}-\frac{1}{R_{0}}\right),&&\text{if }R<R_{0}.\end{aligned}\right.\end{gathered}$
@@ 3.1 A two-dimensional model of early-stage spheroid cutting
Our goal is to understand the effects of a radial incision on early-stage tumour spheroids. Differently from Sec. 2, we first consider a reference configuration where the spheroid is approximated as a cylinder, which is cut radially, with the cylinder’s axis aligned along the $Z$-axis. To simplify the analysis, we assume that deformations along the $Z$-axis are negligible, allowing us to focus solely on deformations in the plane perpendicular to the axis.
@@ 3.1.1 Analytical estimate using simplified kinematics
The transformations introduced in (18) describe the deformation of a half-disk into a disk sector of the same area. By applying this deformation symmetrically to the entire disk, we obtain a “Pac-Man”-like shape, as shown in Fig. 4.
@@
In view of (20) and (21), the total energy (3) becomes
$\mathcal{E}[\boldsymbol{\varphi}_{\alpha}]=\frac{\pi\mu R_{0}^{2}}{4}\left(\alpha+\frac{1}{\alpha}\right)+\gamma R_{0}\left(\frac{1}{\sqrt{\alpha}}+\pi\sqrt{\alpha}\right).$
The energy minima are stationary points of the functional, i.e.
$\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathcal{E}[\boldsymbol{\varphi}_{\alpha}]=0.$ (22)
By solving (22), we get
@@ 3.1.2 Finite element approximation
The implementation is carried out using the FEniCS library. Surface tension is incrementally increased by a step size $\Delta\gamma$. The problem (25) is solved using Newton’s method, with the solution from $\gamma=\gamma_{k}$ serving as the initial guess for the subsequent Newton iteration at $\gamma=\gamma_{k+1}=\gamma_{k}+\Delta\gamma$. The increment $\Delta\gamma$ is automatically adjusted if the Newton method fails to converge or converges in fewer than three iterations. This algorithm is implemented using the library BiFEniCS [Riccobelli_2021].
@@ 3.1.3 Results of the two-dimensional simulations
We observe that for moderate values of the dimensionless parameter $\ell_{c}/R_{o}$ there is a very good agreement between the analytical and the numerical model. When $\ell_{c}/R_{0}>1.5$ the discrepancy between the two models starts to be more relevant. This can be due to the non-homogeneity of the deformations close to the corners, which is particularly evident for large values of $\ell_{c}/R_{0}$. In Fig. 6, we present the current configurations of the cut sphere for various values of the incision length $L$ and of the ratio $\ell_{c}/R_{0}$. A quantitative plot of the different opening length observed as we change the parameters of the model is proposed in Fig. 7, together with a comparison with the experimental results of Guillaume et al. (see [Guillaume_2019]) for a newly created spheroid (i.e. without growth). We observe that surface tension has a strong influence on the opening length (i.e. as we vary $\ell_{c}/R_{0}=\gamma/\mu R_{0}$). Conversely, the incision depth as a smaller influence on the resulting deformations. From the experimental data, we can estimate that $\ell_{c}/R_{0}$ is around $0.15$, even though there is some variance in the experimental results. Given that the shear modulus is approximately $1\,\mathrm{kPa}$ and the radius is around $450\,\mu\mathrm{m}$ [Guillaume_2019], the surface tension acting on the spheroid is estimated to be about $68\,\mathrm{mN/m}$.
@@ 3.2 Simulations of three-dimensional spheroid cutting
In the paper [Guillaume_2019], the authors provides data on the opening lengths observed after cutting spheroids that had been cultured for six days, reporting a doubling of their initial radius. Therefore, in what follows we focus on the case of $R_{T}=2R_{0}$ in the simulations with growth. These spheroids are characterized by an initial cell count of 500, with a final radius $R_{T}\sim 450\mu\mathrm{m}$ at $T=6\,\mathrm{days}$.
@@ 3.2.2 Results of the three-dimensional simulations
The residual stresses accumulated in the spheroid during surface growth nearly double the tumour’s opening compared to simulations without growth, as illustrated in Fig.10 (top). Notably, the simulation results without surface growth (blue lines) are comparable to those shown in Fig.7 for a circular domain, suggesting that the domain geometry plays a minor role in the mechanics of tumour opening post-incision.
A comparison with the experimental measurements of the opening length from Guillaume et al. [Guillaume_2019] is shown in Fig. 10 (bottom). We observe a nearly linear relationship between the incision depth and the measured opening, although the slope of the curves is relatively small. This indicates that incision depth plays a minor role in the overall quantitative opening of the tumour. Instead, surface tension seems to have a more dominant influence. From the experimental results, we estimate the elastocapillary length to be between $0.1R_{T}$ and $0.15R_{T}$. Given that the shear modulus is approximately $1\,\mathrm{kPa}$ and the radius is around $450,\mu\mathrm{m}$ [Guillaume_2019], the surface tension acting on the spheroid is estimated to be in the range of $45\,\mathrm{mN/m}<\gamma<68\,\mathrm{mN/m}$. This range closely aligns with the estimates reported by Riccobelli and Bevilacqua [Riccobelli_2020], who estimated $\gamma\sim 100\,\mathrm{mN/m}$ using a different methodology.
@@ 4 Discussion and concluding remarks
Despite the large amount of literature on the mechanics of solid tumours, even basic aspects of cell proliferation and solid stress development are poorly understood. The theoretical framework of this paper provides a new perspective on these fundamental issues. As discussed in [Riccobelli_2020], we argue that surface tension generated by cell-cell adhesion and active contractility plays a crucial role in determining the stress state and guiding cell proliferation in the early stages of tumour growth.
@@
We have focused our attention on solid tumour spheroids, proposing a mathematical model of tumour surface accretion based on the theory of Zurlo and Truskinovsky [Zurlo_2017, Zurlo_2018, Truskinovsky_2019]. Surface tension elastically deforms newly produced cells on the free surface, resulting into a compression of the cells along the radial direction. This process generates an inhomogeneous prestretch and a geometrically incompatible relaxed state. The pattern of residual stress predicted by the model resembles that observed in real tumours [Stylianopoulos_2012, Ambrosi_2017]. The model predicts a structure of the spheroid composed of an inner core subject to isotropic compression and an outer rim where cells are radially compressed and circumferentially elongated, consistently with experimental observations, see Fig. 5 in [Dolega_2017].
@@
To validate our theoretical framework, we have investigated the deformation of spheroids following radial incisions. In [Guillaume_2019], different opening length are observed depending on the spheroid’s age. Tumour spheroids typically requires two days to self-aggregate [Lee_2019]. After this period, the bulk of the spheroid shows negligible residual stress. Instead, the primary source of stress within the spheroid can be attributed to the tissue surface tension, resulting in a pressure-like stress state in bulk [Lee_2019, Riccobelli_2020]. We have shown that the radial cut generates an opening of the spheroid driven by surface tension, even in the absence of residual stress. We derived a simplified analytical approximation by means of classical universal solutions of nonlinear elasticity, which provides an estimation of tumour opening. However, this model neglects the non-uniform deformations arising from singular stresses at the cut edges, which are caused by the infinite curvature of the free surface in the reference configuration.
@@
To overcome these limitations, we have performed finite element simulations. The results of these simulations are reported in Figs. 5-7. We have shown that, despite its simplicity, our analytical approximation is in good agreement with the results of the numerical simulations. We have also included the effect of growth and the related mechanical stress in older spheroids. The results of the simulations for growing spheroids are reported in Figs. 9-11. The morphology predicted by the numerical simulations strikingly matches with the experimental evidences. The numerical simulations in early and grown spheroids are tested with the quantitative data provided by Guillaume et al. [Guillaume_2019]. Both cases indicate that the surface tension acting on the spheroids is about $60\,\mathrm{mN/m}$, similarly to previous estimates [Riccobelli_2020].
The results of this work highlight the key role that surface tension and elasticity play in the growth of solid tumours. The proposed model correctly captures both the proliferation pattern within the tumour and the development of mechanical stress in the early stages of tumour progression. These key features play a crucial role in the subsequent vascularization of the tumour and possible tissue invasion in cancer. Open problems related to this research are correlated with mechano-transduction phenomena in solid tumours. It is well-known that mechanical stress can inhibit growth, and a mathematical framework to include mechanical feedback in volumetric growth law is quite established. Its extension to surface growth is non-trivial. Other problems involve extending the proposed model to compressible constitutive law. Indeed, the poroelastic nature of the spheroid can induce volumetric deformations of the solid matrix, and adding a slight compressibility would improve the model accuracy. The compressibility of the spheroid can induce discrepancies between the reference and the current configurations, leading to a more complex reconstruction of the surface accretion process, as discussed in [Truskinovsky_2019]. In this respect, also a coupled theory between surface and volumetric growth in biological tissues is needed.

### v2

# Surface tension-driven boundary growth in tumour spheroids

@@
Surface tension-driven boundary growth
@@
D. Riccobelli1, ${}^{1,\,2}\,$ email: davide.riccobelli@sissa.it
(1Mathematics Area, mathLab, SISSA, via Bonomea 265, 34136, Trieste, Italy
2MOX – Dipartimento di Matematica, Politecnico di Milano, Piazza Leonardo da Vinci 32, 20133, Milano, Italy
@@ 1 Introduction
Cancer research frequently relies on surrogates to investigate the dynamics of tumour growth and the effects of drugs, ranging from experiments performed in vivo to in vitro. In the first type of experiments, tumours are implanted and grown in animals, such as mice [Stylianopoulos_2012]. Conversely, in vitro experiments involve the cultivation of tumour cells outside a living organism in the laboratory [Zanoni_2020]. A notable example is provided by multicellular tumour spheroids, which are three-dimensional structures that self-assemble, forming spherical aggregates.
@@
Figure 1: (Left) Sections of tumour spheroids with varying radii. Recently dividing cells are coloured green, while necrotic cells are indicated in blue (adapted from [Browning_2021]). (Right) An incised tumour spheroid two days post-seeding, with an initial cell count of 5000 (adapted from [Guillaume_2019]).
In order to understand how boundary loads drive the inner mechanical state, a rheological description of spheroids is essential. Many authors describe spheroids and cellular aggregates as fluids subject to surface tension [Davis_1997, Foty_1994]. However, solid matter can also exhibit surface tension effects, especially when surface forces dominate over bulk elasticity. This interplay between elasticity and surface tension has given rise to the field of elastocapillarity [style2017elastocapillarity, Bico_2018].
An increasing body of evidence indicates that cellular aggregates exhibit viscoelastic behaviour, with elasticity playing a fundamental role. The mechanical response is strongly influenced by surface tension, particularly in aggregates with small characteristic lengths, such as tumour spheroids. For instance, elasticity and surface tension govern the formation of brain sulci during brain organoid growth [Riccobelli_2020]. Similarly, in [Yadav_2022] the authors analyse the spatial reorganization of tumour cells when a spheroid is locally exposed to laser ablation, inducing a surface tension gradient at the spheroid boundary. Their findings reveal that the resulting cell motion differs from classical Marangoni flow in fluids, underscoring the importance of the spheroid’s elastic properties in the process. Furthermore, in [Ang_2024] the authors investigate elastocapillary effects in glioblastoma cell spheroids embedded in a collagen matrix. Their study demonstrates that matrix contraction and deformation are driven by the interplay of spheroid surface tension, cell-collagen interactions, and poroelastic effects.
Interestingly, similarly to fully developed tumours, spheroids tend to open when subjected to a radial cut [Colin_2018, Guillaume_2019, Kosheleva_2023]. If the spheroid obeyed a fluid-like behaviour governed by surface tension, it would reorganize to minimize surface area, returning to a spherical shape. However, experimental evidence shows that this is not the case. Indeed, radially cut tumour spheroids display a heart-like shape, see Fig. 1. Moreover, even intact tumour spheroids can exhibit both spherical and cylindrical shapes depending on the initial number of cells [Guillaume_2019], suggesting that while surface tension plays a role in the spheroid physics, other forces also influence its shape.
@@
During the initial days, only a thin outer layer of cells experiences circumferential tension, while the core remains in an isotropic stress state [Lee_2019]. This tensional “skin” at the spheroid’s boundary can be understood as an effect of tissue surface tension on the cell aggregate [Riccobelli_2020]. This surface phenomenon arises from the combined effects of passive intercellular cohesion (mediated by E-cadherin) and active cortical contractility (driven by myosin II) [manning2010coaction], with evidence suggesting a greater contribution from the latter [Yadav_2022, Yousafzai_2022]. The coupling of elasticity and surface tension is therefore a key determinant of the mechanical behaviour of cellular aggregates, particularly at the length scales characteristic of multicellular spheroids [Riccobelli_2020, Yadav_2022, Yousafzai_2022, Kosheleva_2023, Ang_2024].
The aim of this work is to construct a mathematical model of tumour spheroid growth able to explain the tumour opening following a radial cut, as reported in the experiments of Guillaume et al. [Guillaume_2019]. In Section 2, we propose the mathematical model of the spheroid, where tumour proliferation is described as a surface accretion mediated by surface tension. The study of spheroid incision is performed in Section 3, where we propose both a simplified analytical prediction of tumour opening and a quantitative analysis through numerical simulations and comparison with experimental results. Finally, the main results are summarized in Section 4, together with some concluding remarks.
2 Surface tension and elasticity in tumour spheroids
@@
2.1 A continuum model of tumour spheroid with surface tension
@@
In order to take into account the large elastic deformations of the tumour, we assume that the cellular aggregate can be described as a hyperelastic material. Although tumour spheroids can accumulate extracellular matrix in their inner regions due to the formation of a necrotic core [nederman1984demonstration], the spheroids used in the experiments by Guillaume et al. [Guillaume_2019] are relatively young and lack a necrotic core. Thus, it is reasonable to assume that the spheroid is homogeneous. Let $\psi_{r}=\psi_{r}(\mathbf{F})$ be the strain energy density. The bulk elastic energy of the body is thus given by
@@
Cellular aggregates are nearly incompressible due to their high content of water. Here, we approximate their behaviour as incompressible, i.e. the deformation must obey the constraint
@@
where $\mu$ is the shear modulus, while :::: denotes the Frobenius scalar product of tensors, i.e. $\mathbf{A}:\mathbf{B}=\operatorname{tr}(\mathbf{A}^{T}\mathbf{B})$.
2.2 Surface growth
While the model proposed in the previous section can describe the elastic behaviour of a spheroid, it is insufficient to capture what occurs as the spheroid grows. Indeed, spheroids left to grow for several days in an environment with large availability of nutrients can generate mechanical stress as a consequence of their growth. The increase in the opening length of cut spheroids grown for a few days [Guillaume_2019] suggests that the growing process enhances the natural opening induced by surface tension by generating a tensile hoop stress in the outer part of the spheroid.
@@ 2.2 Surface growth
Differently from previous models [Mascheroni_2016, Ambrosi_2017, Dolega_2021, Erlich_2023, Walker_2023], here we describe spheroid’s growth as a surface accretion, where new mass is added at the boundary rather than in the bulk. We start by considering a time-dependant reference configuration, i.e. $\Omega_{0}(t)$. In particular, we consider $t\in[0,\,T]$, where $T$ is the final time instant. Let $R_{t}$ be the radius of the (intact) tumour spheroid at time $t$, so that
@@
Such an operation is called material transplant. Generally speaking, the reference configuration may not be stress-free if the transplant map $\mathbf{P}(\boldsymbol{X})$ is not a gradient of a deformation function $\boldsymbol{\varphi}_{\mathbf{P}}(\boldsymbol{X})$, for a detailed discussion see [Epstein_2012, Goriely_2017]. In the following, the explicit dependence of $\mathbf{P}$ on $\boldsymbol{X}$ is omitted unless necessary.
@@
$\Sigma_{RR}=\left\{\begin{aligned} &4\gamma\left(\frac{1}{R_{T}}-\frac{1}{R}\right),&&\text{if }R\geq R_{0},\\ &4\gamma\left(\frac{1}{R_{T}}-\frac{1}{R_{0}}\right),&&\text{if }R<R_{0},\end{aligned}\right.$
$\Sigma_{\Theta\Theta}=\left\{\begin{aligned} &2\gamma\left(\frac{2}{R_{T}}-\frac{1}{R}\right),&&\text{if }R\geq R_{0},\\ &4\gamma\left(\frac{1}{R_{T}}-\frac{1}{R_{0}}\right),&&\text{if }R<R_{0}.\end{aligned}\right.$
@@ 3.1 A two-dimensional model of early-stage spheroid cutting
Our goal is to understand the effects of a radial incision on early-stage tumour spheroids. Differently from Section 2, we first consider a reference configuration where the spheroid is approximated as a cylinder, which is cut radially, with the cylinder’s axis aligned along the $Z$-axis. To simplify the analysis, we assume that deformations along the $Z$-axis are negligible, allowing us to focus solely on deformations in the plane perpendicular to the axis.
@@ 3.1.1 Analytical estimate using simplified kinematics
The transformation introduced in (18) describes the deformation of a half-disk into a disk sector of the same area. By applying this deformation symmetrically to the entire disk, we obtain a “Pac-Man”-like shape, as shown in Fig. 4.
@@
The energy minima are stationary points of the functional $\mathcal{E}[\boldsymbol{\varphi}_{\alpha}]$, i.e.
$\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathcal{E}[\boldsymbol{\varphi}_{\alpha}]=\frac{\mathrm{d}}{\mathrm{d}\alpha}(\mathcal{E}_{\text{el}}[\boldsymbol{\varphi}_{\alpha}]+\mathcal{E}_{\text{s}}[\boldsymbol{\varphi}_{\alpha}])=0,$
which gives
@@ 3.1.2 Finite element approximation
The implementation is carried out using the FEniCS library [langtangen2017solving]. Surface tension is incrementally increased by a step size $\Delta\gamma$. The problem (24) is solved using Newton’s method, with the solution from $\gamma=\gamma_{k}$ serving as the initial guess for the subsequent Newton iteration at $\gamma=\gamma_{k+1}=\gamma_{k}+\Delta\gamma$. The increment $\Delta\gamma$ is automatically adjusted if the Newton method fails to converge or converges in fewer than three iterations. This algorithm is implemented using the library BiFEniCS [Riccobelli_2021].
@@ 3.1.3 Results of the two-dimensional simulations
We observe that for moderate values of the dimensionless parameter $\ell_{c}/R_{o}$ there is a very good agreement between the analytical and the numerical model. When $\ell_{c}/R_{0}>1.5$ the discrepancy between the two models starts to be more relevant. This can be due to the non-homogeneity of the deformations close to the corners. In Fig. 6, we present the current configurations of the cut sphere for various values of the incision length $L$ and of the ratio $\ell_{c}/R_{0}$. A quantitative plot of the different opening length observed as we change the parameters of the model is proposed in Fig. 7, together with a comparison with the experimental results of Guillaume et al. [Guillaume_2019] for a newly created spheroid (i.e. without growth). We observe that surface tension has a strong influence on the opening length (i.e. as we vary $\ell_{c}/R_{0}=\gamma/\mu R_{0}$). Conversely, the incision depth as a smaller influence on the resulting deformations. From the experimental data, we can estimate that $\ell_{c}/R_{0}$ is around $0.15$, even though there is some variance in the experimental results. Given that the shear modulus is approximately $1\,\mathrm{kPa}$ and the radius is around $450\,\mu\mathrm{m}$ [Guillaume_2019], the surface tension acting on the spheroid is estimated to be about $68\,\mathrm{mN/m}$.
@@ 3.2 Simulations of three-dimensional spheroid cutting
In [Guillaume_2019] the authors provides data on the opening lengths observed after cutting spheroids that had been cultured for six days, reporting a doubling of their initial radius. Therefore, in what follows we focus on the case where the final radius is twice the initial radius, namely $R_{T}=2R_{0}$. These spheroids are characterized by an initial cell count of 500, with a final radius $R_{T}\sim 450\mu\mathrm{m}$.
@@ 3.2.2 Results of the three-dimensional simulations
The residual stresses accumulated in the spheroid during surface growth nearly double the tumour’s opening compared to simulations without growth, as illustrated in Fig.10 (top). The simulation results without surface growth (blue lines) are comparable to those shown in Fig.7 for a circular domain, suggesting that the domain geometry plays a minor role in the mechanics of tumour opening post-incision.
A comparison with the experimental measurements of the opening length from Guillaume et al. [Guillaume_2019] is shown in Fig. 10 (bottom). We observe a nearly linear relationship between the incision depth and the measured opening, although the slope of the curves is relatively small. This indicates that incision depth plays a minor role in the overall quantitative opening of the tumour. Instead, surface tension seems to have a more dominant influence. From the experimental results, we estimate the elastocapillary length to be between $0.1R_{T}$ and $0.15R_{T}$. Given that the shear modulus is approximately $1\,\mathrm{kPa}$ and the radius is around $450\,\mu\mathrm{m}$ [Guillaume_2019], the surface tension acting on the spheroid is estimated to be in the range of $45\,\mathrm{mN/m}<\gamma<68\,\mathrm{mN/m}$. This range aligns with the estimates reported by Riccobelli and Bevilacqua [Riccobelli_2020], who estimated $\gamma\sim 100\,\mathrm{mN/m}$ using a different methodology.
@@ 4 Discussion and concluding remarks
Despite the large amount of literature on the mechanics of solid tumours, even basic aspects of cell proliferation and solid stress development are poorly understood. The theoretical framework of this paper provides a new perspective on these fundamental issues. As discussed in [Riccobelli_2020], we argue that surface tension plays a crucial role in determining the stress state and guiding cell proliferation in the early stages of tumour growth.
@@
We have focused our attention on solid tumour spheroids, proposing a mathematical model of surface accretion based on the theory of Zurlo and Truskinovsky [Zurlo_2017, Zurlo_2018, Truskinovsky_2019]. Surface tension elastically deforms newly produced cells on the free surface, resulting in a compression of the cells along the radial direction. This process generates an inhomogeneous prestretch and a geometrically incompatible relaxed state. The pattern of residual stress predicted by the model resembles that observed in real tumours [Stylianopoulos_2012, Ambrosi_2017]. The model predicts a structure of the spheroid composed of an inner core subject to isotropic compression and an outer rim where cells are radially compressed and circumferentially elongated, consistently with experimental observations, see Fig. 5 in [Dolega_2017].
@@
To validate our theoretical framework, we have investigated the deformation of spheroids following radial incisions. In [Guillaume_2019], different opening length are observed depending on the spheroid’s age. Tumour spheroids typically require two days to self-aggregate. After this period, the bulk of the spheroid shows negligible residual stress [Lee_2019]. Instead, the primary source of stress within the spheroid can be attributed to the tissue surface tension, resulting in a pressure-like stress state in bulk [Riccobelli_2020]. We have shown that the radial cut generates an opening of the spheroid driven by surface tension, even in the absence of residual stress. We derived a simplified analytical approximation using classical universal solutions of nonlinear elasticity, which provides an estimation of tumour opening. However, this model neglects the non-uniform deformations arising from singular stresses at the cut edges, which are caused by the infinite curvature of the free surface in the reference configuration.
@@
To overcome these limitations, we have performed finite element simulations. The results of these simulations are reported in Figs. 5-7. We have shown that, despite its simplicity, our analytical approximation is in good agreement with the outcomes of the numerical simulations. We have also included the effect of growth and the related mechanical stress in older spheroids. The results of the simulations for growing spheroids are reported in Figs. 9-11. The morphology predicted by the numerical simulations strikingly matches with the experimental evidence. The numerical simulations in early and grown spheroids are compared with the quantitative data provided by Guillaume et al. [Guillaume_2019]. Both cases indicate that the surface tension acting on the spheroids is about $60\,\mathrm{mN/m}$, similarly to previous estimates [Riccobelli_2020].
The results of this study show the important role played by surface tension and elasticity in the growth of solid tumours, aligning with observations from several studies [Yadav_2022, Yousafzai_2022, Ang_2024]. The proposed model sheds light on the influence of surface tension on tumour growth and elasticity, effectively capturing both the proliferation patterns within the tumour and the development of mechanical stress during the early stages of tumour progression. These fundamental mechanisms are crucial for understanding the subsequent processes of tumour vascularization and potential tissue invasion in cancer.
Open challenges related to this research are closely tied to mechano-transduction phenomena in solid tumours. Mechanical stress is known to inhibit growth, and while a mathematical framework for incorporating mechanical feedback into volumetric growth laws is well-established, its extension to surface growth remains non-trivial. Another key issue involves extending the proposed model to include compressible constitutive laws. Experimental evidence suggests distinct rheologies for the extracellular matrix and cells, with the former being compressible and the latter incompressible [Dolega_2021]. In this respect, the extracellular matrix plays an important role in regulating cell proliferation and motility [Dolega_2021_2, Ang_2024].
While the current model captures the physical mechanisms underlying residual stress generation in tumour spheroids, a more refined model would require the extension to compressible nonlinear elastic and poroelastic materials [Ambrosi_2017, dehghani2020role, Ang_2024]. However, spheroid compressibility can induce discrepancies between the reference and current configurations. This leads to a more complex reconstruction of the surface accretion process, as discussed in [Truskinovsky_2019]. More generally, a coupled theory between surface and volumetric growth and remodelling would be important to accurately model biological tissues.